# On the Exoplanet Yield of Gaia Astrometry

### v1

@@ 1 Introduction
The astrometric method for planet detection requires monitoring the position of a star on the celestial sphere with a precision that has historically been difficult to achieve. To our knowledge, all exoplanet discoveries reported based on ground-based optical astrometry have proven spurious (see, e.g., Hunter, 1943; Holmberg, 1938; Reuyl & Holmberg, 1943; Strand, 1943; Lippincott, 1960; van de Kamp, 1963; Hershey & Lippincott, 1982; Gatewood, 1996; Pravdo & Shaklan, 2009). Very-long-baseline radio interferometry can in principle deliver the necessary precision, but is applicable to few stars, and has yielded only one planet to date (Curiel et al., 2022).
@@
Figure 1: Masses and semi-major axes for known planets and simulated Gaia planets. In both cases, we have restricted the sample to planets with $P_{\mathrm{orb}}$ and $m_{p}$ constrained to within 2020% (i.e., $P_{\mathrm{84th}}/P_{\mathrm{16th}}\,{<}\,1.2$ and $m_{\mathrm{84th}}/m_{\mathrm{16th}}\,{<}\,1.2$). Blue and green points show known planets discovered with the Doppler and transit methods. Red and black points are from our mock Gaia planet catalogs for DR4 and DR5 (see Section 4). For comparison, the values of Jupiter and Saturn are highlighted with gray points labeled “J” and “S.” The Gaia planets are primarily super-Jupiters on multi-AU orbits. About $100$ such planets are currently known; Gaia promises to expand this sample to thousands in DR4 and tens of thousands in DR5.
@@ 2.6 Planet occurrence model
The CLS occurrence rate model does not depend on stellar mass and metallicity, even though giant-planet occurrence is known to vary with those two stellar properties (see Johnson et al. 2010). Assuming that the CLS stars and the relevant Gaia stars have similar metallicity distributions, the CLS model can be used to predict the total number of Gaia planets without accounting explicitly for metallicity. This assumption seems reasonable because the metallicity distribution of the CLS stars (see Figure 4 of Rosenthal et al., 2021) is centered at $0.0$ with a spread from 0.5-0.5 to $0.5$, as observed in the solar neighborhood. However, we must contend with the dependence on stellar mass, because the detectability of planets depends sensitively on stellar mass (see Figures 5 and 6). The occurrence of giant planets as a function of stellar mass has been investigated extensively, at least as far back as Laws et al. (2003). The occurrence rate of giant planets with orbital separations less than a few AU has been found to grow approximately linearly with stellar mass up to  ${\sim}\,2\,M_{\odot}$ (Johnson et al., 2010; Reffert et al., 2015; Jones et al., 2016; Ghezzi et al., 2018; Wolthoff et al., 2022). To incorporate this trend into our planet occurrence model, we allowed the normalization constant $C$ to depend on stellar mass:
@@ 3 Simulated astrometric data
Figure 9: Simulated Gaia DR5 astrometry for an $8\,M_{\mathrm{J}}$ planet on a $4$ AU orbit around a $0.8\,M_{\odot}$ star located $100$ pc away (G= 11G\,{=}\,11). The two panels in the upper right corner show the RA ($\alpha$) and Dec ($\delta$) measurements and the individual along-scan astrometric measurements ($\eta$). For both plots, we removed the best-fit proper motion and overplotted the best-fit planet model (blue). The rest of the figure is a corner plot showing the 2D posterior probability distributions for all model parameters except RA and Dec (which were omitted to keep the figure size manageable). The true values of all parameters (red) were recovered accurately within the uncertainties. This system has SNR1 ${\approx}\,7$; we expect that  ${\sim}\,480$ planets in DR4 and  ${\sim}\,2{,}300$ in DR5 will have SNR1> 7{>}\,7 (see Figure 8).
@@ 3.1 Maximum orbital periods for DR4 and DR5
Figure 10: Accuracy of fitted orbital periods, based on analyses of simulated Gaia astrometric time series data. The top panel shows the recovered orbital period (i.e., the median and $1\sigma$ range) versus the true orbital period for simulated DR4 data. The bottom panel shows the same for simulated DR5 data. For the DR4 data, periods shorter than about $4.0$ years can be accurately recovered. For DR5, the limit is extended to about $9.5$ years.
@@ 4.2 Injection-recovery setup
    The orbital period is well constrained. Specifically, we required $P_{\mathrm{84th}}/P_{\mathrm{16th}}\,{<}\,1.5$, where $P_{\mathrm{xth}}$ is the $x$th percentile of the marginalized posterior probability distribution for the orbital period.
@@
Figure 12: Recovered properties for the 7,5007{,}500 mock Gaia DR4 detections (red) and the 120,000120{,}000 DR5 detections (black). Contours show the $1.0\sigma$ and $2.0\sigma$ ranges in each 2D distribution, and the histograms show the 1D distributions. The plots in the top right corner show how the number of detections depends on the $\Delta\chi^{2}$ threshold and the orbital period precision threshold (see Section 4.3). Most detections will be super-Jupiters on few-AU orbits around GKM-type stars within a few hundred parsecs. The longer timespan of the DR5 data will dramatically increase Gaia’s search volume, resulting in  ${\sim}\,16$ times as many detections as in DR4. Only a subset of the detected planets will have well-constrained masses; we predict that $1{,}900\,{\pm}\,540$ and $38{,}000\,{\pm}\,7{,}300$ planets will have masses and periods measured to within 2020% in DR4 and DR5, respectively (see Section 4.3).
@@ 4.3 Injection-recovery results
Some other biases are more subtle. Even though the simulated planets have random orbital orientations, there is a mild deficit of detected planets with nearly edge-on orbits ($i\,{\approx}\,90^{\circ}$), because astrometry provides less information about their orbits; one dimension is lost to the sky projection (see Gaia Collaboration et al. 2023; Makarov 2025). Additionally, relatively few detected planets have e ${\approx}\,0, e$ven though many simulated planets have low-eccentricity orbits. The reason is that Gaia’s constraints on the eccentricity are generally poor. Since zero is the minimum permissible value of $e$, most planets with e ${\approx}\,0 e$nd up with best-fit eccentricities of $e\,{\approx}\,0.2$ due to noise. This is sometimes referred to as the Lucy & Sweeney (1971) bias.
@@ 5.2 False-positive predictions
We recorded a total of 1,2001{,}200 binary false-positive detections in DR4 and 6,3006{,}300 in DR5. This represents 1313% of the total number of planet candidates (planets and unresolved binaries) in DR4 and 55% in DR5. Although genuine planet detections dominate overall, the fraction of false positives depends strongly on the system parameters. Figure 13 shows the ratio of the number of genuine planet detections to the total number of detections. Relative to the number of planet detections, impostors are more probable when the source has a low apparent mass, a faint apparent magnitude, or a close-orbiting planet. Small orbits are preferred by impostors because the stars undergo lower-amplitude astrometric motion, allowing them to mimic planetary-scale signals without as much fine-tuning of the flux ratio (Equation 40). The preference of impostors for low stellar masses has a similar explanation: at fixed orbital period, lower-mass binaries have smaller orbits. In DR5, the false positive rate for low-mass stars and faint stars will be somewhat lower than in DR4. The DR4 and DR5 false positive rates for close-orbiting planets ( ${\lesssim}\,1$ AU) are expected to be similar.
@@ 5.3 Identifying planet-impostor binaries
Figure 15 compares the colors and magnitudes of the stars hosting detected planets and the planet-impostor binaries. For sources with $1\,{\lesssim}\,G_{BP}\,{-}\,G_{RP}\,{\lesssim}\,2$, unresolved binaries are somewhat brighter than stars with planets, although the difference is modest. Discarding 2020% of the planet impostors using a color-dependent luminosity threshold also removes about 55% of singles with planets. As such, we expect that only a modest fraction of unresolved binaries will be readily identifiable based on their colors and luminosities.
@@ 6.1 Comparison with earlier projections
As mentioned in Section 1, the most comprehensive prior estimates for Gaia’s planet yield were performed by Perryman et al. (2014). Whereas we predict that Gaia will discover $7{,}500\,{\pm}\,2{,}100$ planets in DR4 and $120{,}000\,{\pm}\,22{,}000$ in DR5, Perryman et al. (2014) predicted there would be $21{,}000\,{\pm}\,6{,}000$ in DR4 and $70{,}000\,{\pm}\,20{,}000$ in DR5, respectively. There are many possible reasons for the discrepancies between the predictions, but an especially notable difference is related to the detection criteria. We recorded detections whenever $\Delta\chi^{2}\,{>}\,50$, the orbital period was modestly constrained ($P_{\mathrm{84th}}/P_{\mathrm{16th}}\,{<}\,1.5$), and the planet’s best-fit mass was below $13~M_{\mathrm{J}}$. The uncertainty in our projection reflects the uncertainty in the planet occurrence function. In contrast, Perryman et al. (2014) required detections to satisfy $P_{\mathrm{orb}}\,{<}\,6$ years and $m_{p}\,{<}\,15\,M_{\mathrm{J}}$. The quoted result of $21{,}000\,{\pm}\,6{,}000$ was based on the spread in the number of detections they recorded when considering $\Delta\chi^{2}$ thresholds between $30$ and $50$, which they labeled as “marginal” and “reliable” thresholds, respectively. Because the amplitude of the astrometric signal grows sharply with $P_{\mathrm{orb}}$, their sample is dominated by planets with $P_{\mathrm{orb}}\,{\approx}\,6$ year, which typically have poorly constrained orbits. This also explains why planets in their DR4 simulation extend out to larger orbital separations than planets in our DR4 catalog ($5$ AU versus $3$ AU; see their Figure 2 and our Figure 12). For a more direct comparison, we tried dropping the $P_{\mathrm{84th}}/P_{\mathrm{16th}}\,{<}\,1.5$ condition and instead using a period threshold of $P_{\mathrm{orb}}\,{<}\,6$ years. This increased our predicted number of detections to 16,00016{,}000. Extrapolating our results to allow for planet masses as large as $15\,M_{\mathrm{J}}$ and $\Delta\chi^{2}$ values as small as $40$ (the midpoint of Perryman et al. (2014)’s range) resulted in about 26,00026{,}000 detections. The agreement with Perryman et al. (2014) is good but probably somewhat coincidental, because they adopted astrometric uncertainties that were about 5050% lower for bright stars. The differing stellar populations and planet occurrence models may have compensated for this difference.
@@ 6.3 Summary
  • We predict that Gaia will discover $7{,}500\,{\pm}\,2{,}100$ planets in DR4 and $120{,}000\,{\pm}\,22{,}000$ planets in DR5. The dominant known source of uncertainty for this prediction comes from the uncertainty in the planet occurrence function. The number of candidates that are ultimately identified will depend sensitively on the strictness of the detection criteria. We adopted the following detection criteria: the one-companion model is favored over the no-companion model by $\Delta\chi^{2}\,{>}\,50$, the best-fit planet mass is below $13\,M_{\mathrm{J}}$, and the orbital period is well constrained ($P_{\mathrm{84th}}/P_{\mathrm{16th}}\,{<}\,1.5$). Figure 12 shows how the detection numbers depend on these choices.
@@
  • To facilitate community preparation, we have released our mock Gaia DR4 and DR5 exoplanet catalogs, planet-impostor binary catalogs, and code to reproduce our results. [^12] https://github.com/CalebLammers/GaiaForecasts

### v2

@@ 1 Introduction
The astrometric method for planet detection requires monitoring the position of a star on the celestial sphere with a precision that has historically been difficult to achieve. Many exoplanet discoveries reported based on ground-based optical astrometry have proven spurious (see, e.g., Hunter, 1943; Holmberg, 1938; Reuyl & Holmberg, 1943; Strand, 1943; Lippincott, 1960; van de Kamp, 1963; Hershey & Lippincott, 1982; Gatewood, 1996; Pravdo & Shaklan, 2009), although there remain a few viable candidates (Muterspaugh et al., 2010; Gravity Collaboration et al., 2024). Very-long-baseline radio interferometry can in principle deliver the necessary precision, but is applicable to few stars, and has yielded only one planet to date (Curiel et al., 2022).
@@
Figure 1: Masses and semi-major axes of known planets and simulated Gaia planets. In both cases, we have restricted the sample to planets with $P_{\mathrm{orb}}$ and $m_{p}$ constrained to within 2020% (i.e., $P_{\mathrm{84th}}/P_{\mathrm{16th}}\,{<}\,1.2$ and $m_{\mathrm{84th}}/m_{\mathrm{16th}}\,{<}\,1.2$). Blue and green points are real planets discovered with the Doppler and transit methods. Red and black points are from our mock Gaia planet catalogs for DR4 and DR5 (see Section 4). For comparison, the values of Jupiter and Saturn are highlighted with gray points labeled “J” and “S.” The Gaia planets are primarily super-Jupiters on several-AU orbits. About $100$ such planets are currently known; Gaia promises to expand this sample to thousands in DR4 and tens of thousands in DR5.
@@ 2.6 Planet occurrence model
The CLS occurrence rate model does not depend on stellar mass and metallicity, even though giant-planet occurrence is known to vary with those two stellar properties (see Johnson et al. 2010). Assuming that the CLS stars and the relevant Gaia stars have similar metallicity distributions, the CLS model can be used to predict the total number of Gaia planets without accounting explicitly for metallicity. This assumption seems reasonable because the metallicity distribution of the CLS stars is centered at $0.0$ with a spread from 0.5-0.5 to $0.5$ (see Figure 4 of Rosenthal et al., 2021), as observed in the solar neighborhood. However, we must contend with the dependence on stellar mass, because the detectability of planets depends sensitively on stellar mass (see Figures 5 and 6). The occurrence of giant planets as a function of stellar mass has been investigated extensively, at least as far back as Laws et al. (2003). The occurrence rate of giant planets with orbital separations less than a few AU has been found to grow approximately linearly with stellar mass up to  ${\sim}\,2\,M_{\odot}$ (Johnson et al., 2010; Reffert et al., 2015; Jones et al., 2016; Ghezzi et al., 2018; Wolthoff et al., 2022). To incorporate this trend into our planet occurrence model, we allowed the normalization constant $C$ to depend on stellar mass:
@@ 3 Simulated astrometric data
Figure 9: Simulated Gaia DR5 astrometry for an $8~M_{\mathrm{J}}$ planet on a $4$ AU orbit around a $0.8~M_{\odot}$ star located $100$ pc away (G= 11G\,{=}\,11). The two panels in the upper right corner show the RA ($\alpha$) and Dec ($\delta$) measurements and the individual along-scan astrometric measurements ($\eta$). For both plots, we removed the best-fit proper motion and overplotted the best-fit planet model (blue). The rest of the figure is a corner plot showing the 2D posterior probability distributions for all model parameters except RA and Dec (which were omitted to keep the figure size manageable). Because the distance $r$ and proper motion components $\mu_{\alpha}$ and $\mu_{\delta}$ are tightly constrained, we have plotted their offsets from the true values ($\Delta r$, $\Delta\mu_{\alpha}$, and $\Delta\mu_{\delta}$). The true values of all parameters (red) were recovered accurately within the uncertainties. This system has SNR1 ${\approx}\,7$; we expect that  ${\sim}\,480$ planets in DR4 and  ${\sim}\,2{,}300$ in DR5 will have SNR1> 7{>}\,7 (see Figure 8).
@@ 3.1 Maximum orbital periods for DR4 and DR5
Figure 10: Accuracy of fitted orbital periods, based on analyses of simulated Gaia astrometric time series. The top panel shows the recovered orbital period (i.e., the median and $1\sigma$ range) versus the true orbital period for simulated DR4 data. The bottom panel shows the same for simulated DR5 data. For the DR4 data, periods shorter than about $4.0$ years can be accurately recovered. For DR5, the limit is extended to about $9.5$ years.
@@ 4.2 Injection-recovery setup
    The orbital period is at least moderately constrained. Specifically, we required $P_{\mathrm{84th}}/P_{\mathrm{16th}}\,{<}\,1.5$, where $P_{\mathrm{xth}}$ is the $x$th percentile of the marginalized posterior probability distribution for the orbital period.
@@
Figure 12: Recovered properties for the 7,5007{,}500 mock Gaia DR4 detections (red) and the 120,000120{,}000 DR5 detections (black). Contours show the $1\sigma$ and $2\sigma$ ranges in each 2D distribution, and the histograms show the 1D distributions. The plots in the top right corner show how the number of detections depends on the $\Delta\chi^{2}$ threshold and the orbital period precision threshold (see Section 4.3). Most detections will be super-Jupiters on few-AU orbits around GKM-type stars within a few hundred parsecs. The longer timespan of the DR5 data will dramatically increase Gaia’s search volume, resulting in  ${\sim}\,16$ times as many detections as in DR4. Only a subset of the detected planets will have well-constrained masses; we predict that $1{,}900\,{\pm}\,540$ and $38{,}000\,{\pm}\,7{,}300$ planets will have masses and periods measured to within 2020% in DR4 and DR5, respectively (see Section 4.3).
@@ 4.3 Injection-recovery results
Some other biases are more subtle. Even though the simulated planets have random orbital orientations, there is a mild deficit of detected planets with nearly edge-on orbits ($i\,{\approx}\,90^{\circ}$), because astrometry provides less information about their orbits; one dimension is lost to the sky projection (see Gaia Collaboration et al. 2023; Makarov 2025). Additionally, relatively few detected planets have e ${\approx}\,0, e$ven though many simulated planets have low-eccentricity orbits. The reason is that Gaia’s constraints on the eccentricity are generally poor. Since zero is the minimum permissible value of $e$, most planets with e ${\approx}\,0 e$nd up with best-fit eccentricities of $e\,{\approx}\,0.2$ due to noise. This is sometimes referred to as the Lucy & Sweeney (1971) bias. Host stars are distributed roughly uniformly on the sky, with mild biases imprinted by Gaia’s scanning law.
@@ 5.2 False-positive predictions
We recorded a total of 1,2001{,}200 binary false-positive detections in DR4 and 6,3006{,}300 in DR5. This represents 1313% of the total number of planet candidates (planets and unresolved binaries) in DR4 and 55% in DR5. Although genuine planet detections dominate overall, the fraction of false positives depends strongly on the system parameters. Figure 13 shows the ratio of the number of genuine planet detections to the total number of detections. Relative to the number of planet detections, impostors are more probable when the source has a low apparent mass, a faint apparent magnitude, or a close-orbiting planet. Small orbits are preferred by impostors because this results in lower-amplitude astrometric motion, allowing them to mimic planetary-scale signals without as much fine-tuning of the flux ratio (Equation 40). The preference of impostors for low stellar masses has a similar explanation: at fixed orbital period, lower-mass binaries have smaller orbits. In DR5, the false positive rate for low-mass stars and faint stars will be somewhat lower than in DR4. The DR4 and DR5 false positive rates for close-orbiting planets ( ${\lesssim}\,1$ AU) are expected to be similar.
@@ 5.3 Identifying planet-impostor binaries
Figure 15 compares the colors and magnitudes of the stars hosting detected planets and the planet-impostor binaries. For sources with $1\,{\lesssim}\,G_{BP}\,{-}\,G_{RP}\,{\lesssim}\,2$, unresolved binaries are somewhat brighter than stars with planets, although the difference is modest. Discarding 2020% of the planet impostors using a color-dependent luminosity threshold also removes about 55% of singles with planets. As such, we expect that a fairly small fraction of unresolved binaries will be readily identifiable based on their Gaia colors and luminosities alone.
@@ 6.1 Comparison with earlier projections
As mentioned in Section 1, the most comprehensive prior estimates for Gaia’s planet yield were performed by Perryman et al. (2014). Whereas we predict that Gaia will discover $7{,}500\,{\pm}\,2{,}100$ planets in DR4 and $120{,}000\,{\pm}\,22{,}000$ in DR5, Perryman et al. (2014) predicted there would be $21{,}000\,{\pm}\,6{,}000$ in DR4 and $70{,}000\,{\pm}\,20{,}000$ in DR5, respectively. There are many possible reasons for the discrepancies between the predictions, but an especially notable difference is related to the detection criteria. We recorded detections whenever $\Delta\chi^{2}\,{>}\,50$, the orbital period was at least modestly constrained ($P_{\mathrm{84th}}/P_{\mathrm{16th}}\,{<}\,1.5$), and the planet’s best-fit mass was below $13~M_{\mathrm{J}}$. The uncertainty in our projection reflects the uncertainty in the planet occurrence function. In contrast, Perryman et al. (2014) required detections to satisfy $P_{\mathrm{orb}}\,{<}\,6$ years and $m_{p}\,{<}\,15\,M_{\mathrm{J}}$. The quoted result of $21{,}000\,{\pm}\,6{,}000$ was based on the spread in the number of detections they recorded when considering $\Delta\chi^{2}$ thresholds between $30$ and $50$, which they labeled as “marginal” and “reliable” thresholds, respectively. Because the amplitude of the astrometric signal grows sharply with $P_{\mathrm{orb}}$, their sample is dominated by planets with $P_{\mathrm{orb}}\,{\approx}\,6$ years, which typically have poorly constrained orbits. This also explains why planets in their DR4 simulation extend out to larger orbital separations than planets in our DR4 catalog ($5$ AU versus $3$ AU; see their Figure 2 and our Figure 12). For a more direct comparison, we tried dropping the $P_{\mathrm{84th}}/P_{\mathrm{16th}}\,{<}\,1.5$ condition and instead using a period threshold of $P_{\mathrm{orb}}\,{<}\,6$ years. This increased our predicted number of detections to 16,00016{,}000. Extrapolating our results to allow for planet masses as large as $15\,M_{\mathrm{J}}$ and $\Delta\chi^{2}$ values as small as $40$ (the midpoint of Perryman et al. (2014)’s range) resulted in about 26,00026{,}000 detections. The agreement with Perryman et al. (2014) is good but probably somewhat coincidental, because they adopted astrometric uncertainties that were about 5050% lower for bright stars. The differing stellar populations and planet occurrence models may have compensated for this difference.
@@ 6.3 Summary
  • We predict that Gaia will discover $7{,}500\,{\pm}\,2{,}100$ planets in DR4 and $120{,}000\,{\pm}\,22{,}000$ planets in DR5. The dominant known source of uncertainty for this prediction comes from the uncertainty in the planet occurrence function. The number of candidates that are ultimately identified will depend sensitively on the strictness of the detection criteria. We adopted the following detection criteria: the one-companion model is favored over the no-companion model by $\Delta\chi^{2}\,{>}\,50$, the best-fit planet mass is below $13\,M_{\mathrm{J}}$, and the orbital period is moderately constrained ($P_{\mathrm{84th}}/P_{\mathrm{16th}}\,{<}\,1.5$). Figure 12 shows how the detection numbers depend on these choices.
@@
  • To facilitate community preparation, we have released our mock Gaia DR4 and DR5 exoplanet catalogs, planet-impostor binary catalogs, and code to reproduce our results. [^12] https://github.com/CalebLammers/GaiaForecasts. A frozen version is available on Zenodo: 10.5281/zenodo.17568375 (catalog doi: 10.17909/T9XG63).